OpenStudy (anonymous):

Simplify the Product. √5(√15-3)

OpenStudy (anonymous):

\[\sqrt{5}(\sqrt{15}-3)\]first you want to multiply square root of 5 by what is inside the bracket... \[= \sqrt{75} - 3\sqrt{5}\]Next you want to see if you can break down the square toot of 75... \[\sqrt{75}= \sqrt{25 * 3}\] you can further break down the square root of 25 so... \[\sqrt{75} = 5\sqrt{3}\] your final answer is \[5\sqrt{3} - 3\sqrt{5}\]